# Geometry Aware Operator Transformer As An Efficient And Accurate Neural Surrogate For PDEs On Arbitrary Domains

### v2

@@ 1 Introduction
Compared to baselines results taken from the leaderboard of the DrivAerNet++ challenge [8], we see that GAOT significantly improves on the state-of-the-art (see also Fig. 4). This improvement is most visible in the MAE for wall shear stress where GAOT is ca. $30\%$ more accurate than the second-best model (TripNet), which currently sits atop the leaderboard. We recall that GAOT’s decoder endows it with neural field properties. We showcase it for the DrivAerNet++ dataset by training GAOT on a randomly selected set of less than $10\%$ of the total input points (per batch) and then testing on the original car surface point cloud by querying the desired points through GAOT’s decoder. Although not as accurate as training GAOT with full input, we observe from Tab. 3 that this neural field version of GAOT has comparable accuracy to some of the baselines which have been trained with $10$x more input points, further demonstrating the flexibility and accuracy of GAOT.
@@
We highlight GAOT’s flexibility with respect to the point distributions that it can handle by testing it on PDEs with regular grid inputs, as suggested in [39]. To this end, we considered 7 additional datasets and present the test errors in SM Sec. E.3. to find that GAOT is highly accurate even on regular grids and is either more accurate or comparable to the highly expressive GNN-based RIGNO, while being more accurate than widely used neural operators such as FNO and CNO. A key requirement in operator learning [22, 4] is the ability of the model to generalize (at test time) to input and output resolutions that are different from the training resolution. As GAOT can be readily evaluated at any query point, we showcase this aspect of GAOT in SM Sec. E.5. by plotting the test errors for a sequence of resolutions, different from the training resolution, for the Poisson-Gauss benchmark, to find that GAOT generalizes very well in both the sub- and super-resolution settings, even to grids with $10$x more input points than the training resolution. Another test of the generalizability of a model is its ability to perform well out-of-distribution, either zero-shot or when it is fine-tuned with a few in-distribution samples for the new learning task. To test this aspect, we consider the datasets for compressible flow past bluff bodies and train a GAOT model on a set of bluff body shapes and then test it on shapes that were not in the training set. Then, the model is fine-tuned with a few task-specific samples and the results are shown in Fig. 3 (c). We observe that our model performs very well in a few-shot transfer learning scenario, with the fine-tuned model providing an almost order of magnitude gain in accuracy over the model, trained from scratch. Finally, in SM Sec. E.4, we demonstrate that GAOT scales with both model and dataset size, with scaling with respect to dataset size, also illustrated in Fig. 3 (c).
Ablations.
@@
Summarizing the extensive ablation studies of SM Sec.F, we observed that i) the MAGNO encoder/decoder is clearly superior to message-passing based encoders/decoders, ii) choosing a regular equispaced latent point cloud performs significantly better than either downsampling on the original point cloud or using a projected low-dimensional regular grid, iii) a time-derivative marching strategy, i.e, setting $\gamma=1,\delta=\tau$ in (7) is superior to other choices of $\gamma,\delta$, iv) using a statistical geometric embedding performs significantly better than either not using additional geometric information or using a pointnet to process geometric information and v) incorporating multiscale features in the MAGNO encoder/decoder provides a significant gain in accuracy when compared to using just a single scale GNO encoder/decoder as in GINO. A posteriori, these observations justify the choices that we have made in designing GAOT and selecting the relevant model components, while also revealing how these innovative features underpin GAOT’s overall performance.
@@ 4 Discussion
Summary. We present GAOT, a new neural operator for learning PDE solutions on arbitrary domains. It is based on a novel multiscale GNO encoder/decoder, combined with geometric embeddings that convey statistical information about the local domain geometry, and a (vision) transformer based processor. The model is designed to handle any point cloud input and provide the output at any query point in the underlying domain. Several innovative strategies have been used to make the implementation of GAOT computationally efficient and scalable. We test GAOT on a large number of challenging datasets for a variety of time-dependent and time-independent PDEs over diverse two-dimensional domain geometries to find that GAOT is significantly more accurate, robust and computationally efficient in terms of training throughput and inference latency, over a large set of baselines. We further demonstrate the potential of GAOT by presenting its SOTA performance on a large-scale three-dimensional dataset of industrial simulations. These results demonstrate that GAOT can be a powerful and scalable neural operator with wide-spread applications.
@@
Related Work. As discussed before, there are 3 broad classes of models for learning PDEs on arbitrary domains namely i) end-to-end message-passing based frameworks exemplified here with RIGNO, which significantly improves upon models such as (multiscale) MeshGraphNets [41, 15] ii) Transformer based frameworks such as Transolver [50], GNOT [18] and UPT [1] and iii) frameworks, based on GNO encoders/decoders and FNO processors as in GINO [29]. GAOT differs from all these approaches by a) not using graph-based message passing, b) only employing transformers in the processor c) using a transformer, rather than FNO as a processor and significantly augmenting the GNO encoder/decoder by multiscale features, attention based-quadrature and geometry embeddings. It is precisely these choices, along with a highly efficient implementation, that allows GAOT to significantly surpass GINO, RIGNO, and Transsolver in both accuracy and efficiency.
@@
Limitations and Extensions. GAOT’s excellent scalability and ability to generalize very well in a transfer learning setting (Fig. 3(c)) showcase its potential to serve as the backbone of foundation models for PDEs, extending models such as Poseidon [20] and DPOT [17] to arbitrary domains. Further evaluation of GAOT, beyond the DrivAerNet++ dataset, on three-dimensional datasets is also envisaged in future work. Physics-informed loss functions can be added to GAOT to enable it to act as a physics-informed neural operator as in [30]. We also plan to apply GAOT to downstream tasks such as UQ [34], inverse problems, [37] and PDE constrained optimization [35] to further test its abilities. Finally, our main aim in this paper has been to propose GAOT and test it empirically. We have not yet explored rigorous theoretical properties such as universal approximation results [25, 23]. This limitation will also be addressed in future work.
@@ Acknowledgements
The contribution of S. Mishra to this work was supported in part by the DOE SEA-CROGS project (DE-SC-0023191).
@@ B.3.1 Integration of Geometry Embeddings in MAGNO
[…]
@@ B.4 Processor
B.5 Training Details.
@@ B.5.2 Data Normalization
B.5.3 General Training Setup
This section provides an overview of our training hyperparameters. Unless otherwise noted, all experiments follow these settings. Table B.1 summarizes the primary hyperparameters and training schedules. In particular, we distinguish between time-dependent and time-independent PDE tasks in terms of epoch count, and highlight the differences in hardware usage for the DrivAerNet++ dataset. All models except DrivAerNet++ run on a single GeForce RTX 4090 GPU. For the DrivAerNet++ dataset, we use 4 GeForce A100 (40GB) GPUs in data parallel mode, and each GPU holds a batch size of 1. For the scheduler, we warm up to mitigate instability at early epochs, then adopt a cosine schedule for gradual decay, and finalize with a step-based drop for fine-tuning the last epoch range.
@@ B.5.3 General Training Setup
Hardware Single-GPU: All models (except DrivaerNet++) are trained on a single GeForce RTX 4090 with batch size $=64$. Four-GPU: For the DrivaerNet++ dataset, we use four GeForce A100 (40GB) GPUs, each with batch size $=1$.
@@
Epochs Time-Dependent PDEs: $500$ epochs Time-Independent PDEs: $1000$ epochs except the DrivaerNet++ dataset, which is trained on $200$ epochs.
@@
B.5.4 GAOT Model Configuration
@@
FFN 1024 4×4$\times$4 × hidden size (THS) for the feedforward layer.
Table B.2 outlines the default configuration of our GAOT framework. This includes the MAGNO used in both the encoder and decoder stages, as well as the Transformer-based global processor. MAGNO converts node features into geometry-aware tokens (encoder) and reconstructs continuous fields (decoder). By default, the coordinates will be rescaled in the domain $[-1,1]^{d}$, and we use a single aggregation radius $0.033$ for adequate coverage. If multiscale is enabled, we adopt radii $\{0.022,0.033,0.044\}$. The default geometric embedding method is local Statistical Embedding (e.g. as SM Sec. B.3), and will typically be implemented for unstructured datasets. The Transformer processes geometry-aware tokens globally via multi-head self-attention. We set the hidden dimension to $256$ with a $1024$-dim feed-forward layer, residual connections, RMSNorm, and RoPE for positional embeddings. By adjusting the patch size and the number of tokens, we can trade off computational cost and model resolution.
B.6 Inference
@@ C.1 UPT
where $\mathcal{E}$ (Encoder) compresses $k$ input points—coming from an arbitrary Eulerian mesh or Lagrangian particle cloud—into a fixed set of $n_{\text{latent}}$ tokens. It first embeds the features and coordinates through a radius-graph message-passing layer that aggregates information into $n_{s}$ supernodes, and finally employs transformer and perceiver pooling blocks to obtain the latent representation $z_{t}\in\mathbb{R}^{n_{\text{latent}}\times h}$
@@
[…]
@@ Appendix D Datasets
In this work, we test GAOT on 24 benchmarks for both time-independent and time-dependent PDEs of various types, ranging from regular grids to random point clouds to highly unstructured adapted grids. The time-dependent and Poisson-Gauss dataset are sourced from [20] and [39], respectively. A static elasticity dataset is from  [26]. We have generated five additional challenging datasets: a Poisson-C-Sines dataset exhibiting multiscale properties, and four datasets for compressible fluid dynamics with highly unstructured adapted grids. Detailed information regarding these datasets can be found in the Tab D.1.
Table D.1: Overview of the datasets used in this work. Datasets listed above line are time-independent, while those below are time-dependent. Geometry variation (GeoVar) describes whether all data samples share the same geometry. Characteristic briefly describes each dataset’s geometry or PDE setup. The PDE Type column indicates the corresponding class. Visualization (Vis.) provides references to visual examples; for time-dependent datasets, this may include visualizations for both unstructured partially ones and original regular grid ones. Datasets marked with are newly proposed in this work.
@@
NS-Gauss F Gaussian vorticity IC INS G.10, G.18
NS-PwC F Piecewise const. IC INS G.11, G.19
NS-SL F Shear layer IC INS G.12, G.20
NS-SVS F Sinusoidal vortex sheet IC INS G.13, G.21
CE-Gauss F Gaussian vorticity IC CE G.14, G.22
CE-RP F 4-quadrant RP CE G.15, G.23
Wave-Layer F Layered wave medium WE G.16, G.24
Wave-C-Sines F Circular domain with sines IC WE G.17
We focus on the following PDE Types under various initial/boundary conditions and domain geometries:
@@
[…]
@@
$-\,\Delta u\;=\;f,\quad\text{in}\;\,(0,1)^{2},$ (D.2)
with homogeneous Dirichlet boundary conditions. The dataset related to poisson equation use either sinusoidal or Gaussian-like source terms on square or circular domains (see Table D.1).
@@
where $\mathbf{v}$ is the velocity field, $p$ is the pressure, and viscosity is $\nu$. It assumes periodic boundary conditions and sample various initial conditions (e.g., Gaussian, piecewise-constant, sinusoidal vortex sheets).
@@
[…]

### v3

@@ 1 Introduction
Given the high accuracy and excellent computational efficiency and scalability of GAOT, we showcase its abilities further on three challenging three-dimensional large-scale benchmarks for industrial simulations. We start with the DrivAerNet++ dataset of dvnet++ . In this benchmark, the data consists of high-fidelity CFD simulations across 88K different car shapes which span the entire range of conventional car design. The underlying task is to learn steady-state surface fields (See Fig. 4) such as the pressure and wall shear stress, given the input car shape and flow conditions. The data has approximately 500500K points per shape, making the overall training extremely compute intensive. Thus, only scalable models can currently process this learning task. We test GAOT on this challenging 3D benchmark and report the RMSE and MAE test errors for the pressure and wall shear stress in Tab. 3. Compared to baselines results taken from the leaderboard of the DrivAerNet++ challenge tripnet , we see that GAOT significantly improves on the state-of-the-art (see also Fig. 4). This improvement is most visible in the MAE for wall shear stress where GAOT is ca. $30\%$ more accurate than the second-best model (TripNet), which currently sits atop the leaderboard for wall shear stress predictions. We recall that GAOT’s decoder endows it with neural field properties. We showcase it for the DrivAerNet++ dataset by training GAOT on a randomly selected set of less than $10\%$ of the total input points (per batch) and then testing on the original car surface point cloud by querying the desired points through GAOT’s decoder. Although not as accurate as training GAOT with full input, we observe from Tab. 3 that this neural field version of GAOT has comparable accuracy to some of the baselines which have been trained with 1010x more input points, further demonstrating the flexibility and accuracy of GAOT. Further assessments of the neural field property of GAOT are provided in SM E.8.
Next, we consider the very challenging DrivAerML dataset of ashton2024drivaer (see Fig. 5), where the learning task is exactly the same as in DrivAernet++, i.e., predicting surface fields on the car, given its shape as the input. However, unlike DrivAernet++ which was based on coarse RANS simulations, DrivAerML’s ground truth is based on highly accurate LES simulations. This enables the incorporation of much more fine-scale physical effects in this dataset. The learning problem becomes harder, not just in terms of the challenging underlying physics, but also the fact that the number of points on the car surface is now 9M, instead of 500K for DrivAernet++. Thus, only highly scalable models can deal with this extreme resolution. Consequently, we are only able to test GAOT and GINO for this dataset and report the results in Table 4 to observe that GAOT is significantly (almost twice on wall shear-stress) as accurate as GINO on this dataset.
Table 4: Comparison of GAOT and GINO across two benchmarks with MSE ($\times 10^{-2}$) and Mean AE ($\times 10^{-1}$). Cp: Pressure Coefficient, WSS: Wall Shear Stress, P: Pressure, Cf: Skin Friction Coefficient. DML: DrivaerML dataset, CRM: NASA CRM dataset.
Model Cp (DML) WSS (DML) P (CRM) Cf (CRM)
MSE Mean AE MSE Mean AE MSE Mean AE MSE Mean AE
GAOT 5.1729 1.2352 16.9818 2.1640 7.7170 1.6014 16.1091 2.2305
GINO 8.8124 1.5238 28.4832 2.7330 10.5688 1.7450 21.1789 2.4240
Figure 5: Comparison of predicted and ground-truth (GT) results for the pressure on the test sample of DrivAerML and NASA-CRM.
Finally, we consider an industrial-scale dataset, recently proposed in bekemeyer2025introduction , where the learning task (see Fig. 5), is to predict surface pressure and the skin friction coefficient, given the shape of a full aircraft. The ground truth is generated with RANS simulations using a Spalart-Allmaras turbulence model, and the results with GAOT and GINO are reported in Table 4, showing that GAOT significantly outperforms GINO on this large-scale industrial dataset.
@@
We highlight GAOT’s flexibility with respect to the point distributions that it can handle by testing it on PDEs with regular grid inputs, as suggested in rigno . To this end, we considered 7 additional datasets and present the test errors in SM Sec. E.5. to find that GAOT is highly accurate even on regular grids and is either more accurate or comparable to the highly expressive GNN-based RIGNO, while being more accurate than widely used neural operators such as FNO and CNO. A key requirement in operator learning kovachki2023neural ; bartolucci2023representation is the ability of the model to generalize (at test time) to input and output resolutions that are different from the training resolution. As GAOT can be readily evaluated at any query point, we showcase this aspect of GAOT in SM Sec. E.7. by plotting the test errors for a sequence of resolutions, different from the training resolution, for the Poisson-Gauss benchmark, to find that GAOT generalizes very well in both the sub- and super-resolution settings, even to grids with 1010x more input points than the training resolution. Another test of the generalizability of a model is its ability to perform well out-of-distribution, either zero-shot or when it is fine-tuned with a few in-distribution samples for the new learning task. To test this aspect, we consider the datasets for compressible flow past bluff bodies and train a GAOT model on a set of bluff body shapes and then test it on shapes that were not in the training set. Then, the model is fine-tuned with a few task-specific samples and the results are shown in Fig. 3 (c). We observe that our model performs very well in a few-shot transfer learning scenario, with the fine-tuned model providing an almost order of magnitude gain in accuracy over the model, trained from scratch. Finally, in SM Sec. E.6, we demonstrate that GAOT scales with both model and dataset size, with scaling with respect to dataset size, also illustrated in Fig. 3 (c)
Why does GAOT work so well ?
@@
To answer this question, we have performed extensive ablation studies in SM Sec.F to observe that i) the MAGNO encoder/decoder is clearly superior to message-passing based encoders/decoders, ii) choosing a regular equispaced latent point cloud performs significantly better than either downsampling on the original point cloud or using a projected low-dimensional regular grid, iii) GAOT is highly robust to the size of its latent grid, iv) a time-derivative marching strategy, i.e, setting $\gamma=1,\delta=\tau$ in (7) is superior to other choices of $\gamma,\delta$, v) using a statistical geometric embedding performs significantly better than either not using additional geometric information or using a pointnet to process geometric information vi) incorporating multiscale features in the MAGNO encoder/decoder provides a significant gain in accuracy when compared to using just a single scale GNO encoder/decoder as in GINO and vii) The power of GAOT does not just stem from its VIT processor, but also from its MAGNO encoder/decoder (SM E.5), acting in tandem. These results justify the choices that we have made in designing GAOT and selecting the relevant model components, while also revealing how these innovative features underpin GAOT’s accuracy.
However, as argued before, this accuracy might come at the price of computational inefficiency. But, as we have demonstrated above, GAOT is also the most efficient model and does not have to pay the accuracy-efficiency trade-off. The reasons behind this boil down to the tricks used in efficiently implementing GAOT, which are discussed at length in SM E.2 and E.1.
@@ 4 Discussion
Summary. We present GAOT, a new neural operator for learning PDE solutions on arbitrary domains. It is based on a novel multiscale GNO encoder/decoder, combined with geometric embeddings that convey statistical information about the local domain geometry, and a (vision) transformer based processor. The model is designed to handle any point cloud input and provide the output at any query point in the underlying domain. Several innovative strategies have been used to make the implementation of GAOT computationally efficient and scalable. We test GAOT on a large number of challenging datasets for a variety of time-dependent and time-independent PDEs over diverse two-dimensional domain geometries to find that GAOT is significantly more accurate, robust and computationally efficient in terms of training throughput and inference latency, over a large set of baselines. We further demonstrate the potential of GAOT by presenting its SOTA performance on three large-scale three-dimensional datasets of industrial simulations in the automobile and aerospace sectors. These results demonstrate that GAOT can be a powerful and scalable neural operator with wide-spread applications. They also showcase the main advantage with an efficient and accurate neural operator such as GAOT, i.e, its inference time is many orders of magnitude faster than the runtime of a classical numerical PDE solver. We quantify this speedup in SM E.11 to find that GAOT can be anywhere between 4 to 9 orders of magnitude faster to run than classical PDE solvers, while retaining accuracy.
@@
Related Work. As discussed before, there are 3 broad classes of models for learning PDEs on arbitrary domains namely i) end-to-end message-passing based frameworks exemplified here with RIGNO, which significantly improves upon models such as (multiscale) MeshGraphNets MGN ; MSMGN ii) Transformer based frameworks such as Transolver wu2024transolver , GNOT hao2023gnot and UPT UPT and iii) frameworks, based on GNO encoders/decoders and FNO processors as in GINO li2023geometryinformed . GAOT differs from all these approaches by a) not using graph-based message passing, b) only employing transformers in the processor c) using a transformer, rather than FNO as a processor and significantly augmenting the GNO encoder/decoder by multiscale features, attention based-quadrature and geometry embeddings. It is precisely these choices, along with a highly efficient implementation, that allows GAOT to significantly surpass GINO, RIGNO, and Transolver in both accuracy and efficiency.
@@
Limitations and Extensions. GAOT’s excellent scalability and ability to generalize very well in a transfer learning setting (Fig. 3(c)) showcase its potential to serve as the backbone of foundation models for PDEs, extending models such as Poseidon herde2024poseidon and DPOT hao2024dpot to arbitrary domains. Physics-informed loss functions can be added to GAOT to enable it to act as a physics-informed neural operator as in li2024physics . We also plan to apply GAOT to downstream tasks such as UQ LMR1 , inverse problems, NIO and PDE constrained optimization ISMO to further test its abilities. Finally, theoretical results for GAOT, such as universal approximation, will be considered in future work.
@@ Acknowledgements
The contribution of Siddhartha Mishra to this work was supported in part by the DOE SEA-CROGS project (DE-SC-0023191).
@@ B.3.1 Integration of Geometry Embeddings in MAGNO
    where $\mathbin{\|}$ denotes concatenation. And $\mathrm{MLP}_{\mathrm{fuse}}^{m}$ is shared across scales.
@@ B.4 Processor
B.5 Graph Building Tricks
For 2D datasets, we construct neighborhoods for each latent token using a fixed-radius rule, following the Ref. [28]. This works well because local density variations are moderate and the total edge count remains manageable. However, this strategy is not scalable for large 3D industrial datasets, where meshes could be highly adaptive. A single global radius $r$ will lead to a sharp trade-off: if $r$ is large, the number of edges explodes in dense regions and quickly exceeds GPU memory; if $r$ is small, the Overlap of the local neighborhoods fails to cover the whole physical domain and induces information loss. To simultaneously control edge count and guarantee the whole domain coverage, we introduce a bidirectional graph-building strategy that merges two complementary graphs:
  1. 1.
    Radius graph centered at latent tokens. For each latent token $y_{\ell}$, we connect all physical points $x_{j}$ within a radius $r$, producing edges that encode local aggregation around latent tokens. This yields good locality, but may not be able to cover the physical regions if $r$ is too small.
  2. 2.
    KNN graph centered at physical points. For each physical grid point (node) $x_{j}$, we connect it to its $k$ nearest latent tokens, adding at least one attachment per physical node. This forms a lightweight backbone that ensures every physical location is represented in the latent space, even where radius neighborhoods are empty.
We then merge the two edge sets (remove the repeated pairs). The resulting bidirectional graph not only ensures all the physical points are considered, but control the total edge count through the tunable hyperparameter $k$ (typically $k\in[1,4]$) and radius $r$. In practice, we use the encoder direction physical $\to$ latent and the decoder direction latent $\to$ physical; the same bidirectional recipe applies to both by swapping source/target roles, shown in Fig. B.3. Although the bidirectional build is more expensive than a pure-radius build, we precompute graphs offline and cache them, so the extra cost does not affect training throughput.
On all 2D benchmarks, we retain the pure-radius construction. On 3D industrial-scale datasets, we switch to the bidirectional construction to avoid radius-induced edge explosion while preserving coverage. To further control memory and improve robustness, we also apply edge masking during training.
[…]
Figure B.3: Neighbor strategy dispatcher used in GAOT. For 2D datasets, a radius-based graph is adopted; for large 3D datasets, the bidirectional strategy (radius + KNN) is used to ensure domain coverage without edge explosion.
B.6 Training Details.
@@ B.6.2 Data Normalization
B.6.3 General Training Setup
This section provides an overview of our training hyperparameters. Unless otherwise noted, all experiments follow these settings. Table B.1 summarizes the primary hyperparameters and training schedules. In particular, we distinguish between time-dependent and time-independent PDE tasks in terms of epoch count and highlight the differences in hardware usage for the industrial 3D dataset. All models except industrial 3D dataset run on a single GeForce RTX 4090 GPU. For the DrivAerNet++ dataset, we use 4 GeForce A100 GPUs in data parallel mode for 50 epochs, and each GPU holds a batch size of 1. For DrivAerML and NASA-CRM dataset, we use 1 GeForce A100 GPUs for 200 epochs. For the scheduler, we warm up to mitigate instability at early epochs, then adopt a cosine schedule for gradual decay, and finalize with a step-based drop for fine-tuning the last epoch range.
@@ B.6.3 General Training Setup
Hardware Single-GPU: All models for 2D dataset are trained on a single GeForce RTX 4090 with batch size =64=64. Single-GPU: Models for DrivAerML and NASA-CRM are trained on single GeForce A100 GPUs with batch size =1=1. Four-GPU: For the DrivaerNet++ dataset, we use four GeForce A100 GPUs, each with batch size =1=1.
@@
Epochs 2D Time-Dependent PDEs: $500$ epochs 2D Time-Independent PDEs: $1000$ epochs 3D DrivAerNet++: $50$ epochs. 3D NASA-CRM and DrivAerML: $200$ epochs.
@@
B.6.4 GAOT Model Configuration
@@
FFN 1024 $4\times$ hidden size (THS) for the feedforward layer.
Table B.2 outlines the default configuration of the GAOT framework. This includes the MAGNO used in both the encoder and decoder stages, as well as the Transformer-based global processor. MAGNO converts node features into geometry-aware tokens (encoder) and reconstructs continuous fields (decoder). By default, the coordinates will be rescaled in the domain $1,1]^{d}$, and we use a single aggregation radius $0.033$ for adequate coverage. If multiscale is enabled, we adopt radii $\{0.022,0.033,0.044\}$. The default geometric embedding method is local Statistical Embedding (e.g. as SM Sec. B.3), and will typically be implemented for unstructured datasets. The Transformer processes geometry-aware tokens globally via multi-head self-attention. We set the hidden dimension to $256$ with a $1024$-dim feed-forward layer, residual connections, RMSNorm, and RoPE for positional embeddings. By adjusting the patch size and the number of tokens, we can trade off computational cost and model resolution, discussed in SM Sec. F.2.
The configuration slightly differs for the industrial 3D datasets to accommodate large-scale adaptive meshes. We employ the following number of latent tokens for each dataset:
  • NASA-CRM: [64, 64, 32],
  • DrivAerML: [64, 32, 32],
  • DrivAerNet++: [64, 32, 32].
In these setups, the graph construction follows the bidirectional strategy described in Appendix B.5, using a single-scale radius of $0.033$ and a KNN size of $k=1$. The Transformer processor adopts a patch size of $2$ and consists of $10$ layers. All other hyperparameters remain consistent with the default configuration in Table B.2.
B.7 Inference
@@ C.1 UPT
where $\mathcal{E}$ (Encoder) compresses $k$ input points—coming from an arbitrary Eulerian mesh or Lagrangian particle cloud—into a fixed set of $n_{\text{latent}}$ tokens. It first embeds the features and coordinates through a radius-graph message-passing layer that aggregates information into $n_{s}$ supernodes, and finally employs transformer and perceiver pooling blocks to obtain the latent representation $z_{t}\in\mathbb{R}^{n_{\text{latent}}\times h}$ .
@@
In the setting of time-independent problems, we bypass the latent roll-out stage, and adopt a lightweight configuration in our experiments. Specifically, we use latent tokens = $64$ and embedding dimensions =64=64, which results in a model size of 0.74MM. A large variant with $256$ latent tokens and $192$ embedding dimension as setup in [1] was found to suffer from optimization difficulties and was not adopted in our baseline results. The same optimizer setup as GAOT is used here.
@@ Appendix D Datasets
In this work, we test GAOT on 28 benchmarks for both time-independent and time-dependent PDEs of various types, ranging from regular grids to random point clouds to highly unstructured adapted grids. The time-dependent and Poisson-Gauss datasets are sourced from [21] and [39], respectively. The static elasticity dataset is from  [25]. The DrivAerNet++ and DrivAerML datasets are taken from [13] and [2], respectively, where the objective is to predict the surface pressure and wall shear stress. The NASA-CRM dataset is obtained from [6], where we similarly predict the pressure field and surface friction coefficient on the NASA Common Research Model (CRM) surface. For 3D car and wing aerodynamics, the full-resolution surface field is directly predicted without any downsampling. Furthermore, we have also generated five additional challenging datasets: a Poisson-C-Sines dataset exhibiting multiscale properties, and four datasets for compressible fluid dynamics with highly unstructured adapted grids. Detailed information regarding these datasets can be found in the Tab D.1.
Table D.1: Overview of the datasets used in this work. Datasets listed above the line are time-independent, while those below are time-dependent. Geometry variation (GeoVar) describes whether every data sample in the dataset has a different geometry. Characteristics briefly describe each dataset’s geometry or PDE setup. The PDE Type column indicates the corresponding class. Visualization (Vis.) provides references to visual examples; for time-dependent datasets, this may include visualizations for both unstructured partially ones and original regular grid ones. Datasets marked with are newly proposed in this work.
@@
DrivAerML(p) T Surface pressure coefficient INS G.10
DrivAerML(wss) T Surface wall shear stress INS G.11
NASA-CRM (p) T Surface pressure INS G.12
NASA-CRM (sfc) T Surface friction coefficient INS G.13
NS-Gauss F Gaussian vorticity IC INS G.14, G.22
NS-PwC F Piecewise const. IC INS G.15, G.23
NS-SL F Shear layer IC INS G.16, G.24
NS-SVS F Sinusoidal vortex sheet IC INS G.17, G.25
CE-Gauss F Gaussian vorticity IC CE G.18, G.26
CE-RP F 4-quadrant RP CE G.19, G.27
Wave-Layer F Layered wave medium WE G.20, G.28
Wave-C-Sines F Circular domain with sines IC WE G.21
We focus on the following PDE Types simulated for various initial/boundary conditions and domain geometries:
@@
$\rho^{s}\,\frac{\partial^{2}\mathbf{u}}{\partial t^{2}}\;+\;\nabla\cdot\bm{\sigma}\;=\;0,$ (D.1)
@@
$\Delta u\;=\;f,\quad\text{in}\;\,(0,1)^{2},$ (D.2)
with homogeneous Dirichlet boundary conditions. The dataset related to the Poisson equation uses either sinusoidal or Gaussian-like source terms on square or circular domains (see Table D.1).
@@
where $\mathbf{v}$ is the velocity field, $p$ is the pressure, and viscosity is $\nu$. It assumes periodic boundary conditions and samples various initial conditions (e.g., Gaussian, piecewise-constant, sinusoidal vortex sheets).
@@
$\partial_{t}\mathbf{u}\;+\;\nabla\cdot\mathbf{F}\;=\;0,\qquad\mathbf{u}=(\rho,\,\rho\mathbf{v},\,E)^{\top},\quad E=\tfrac{1}{2}\rho\|\mathbf{v}\|^{2}+\tfrac{p}{\gamma-1},$ (D.3)